# E2ENet: Dynamic Sparse Feature Fusion for Accurate and Efficient 3D Medical Image Segmentation

### v1

@@ 3.2 Dynamic Sparse Feature Fusion
19:              $IS\leftarrow$ importance score ($L1$-norm of corresponding kernel) for activated each feature connection
@@
Figure 4: Illustration of our Dynamic Sparse Feature Fusion (DSFF) mechanism. The fusion operation starts from sparse feature connections and allows the connectivity to be evolved after training for $\Delta T$ epochs. During each evolution stage, a fraction of kernels with smaller $L_{1}$-norms will be zeroed out (red dotted line), while the same fraction of other inactivated connections will be reactivated randomly, keeping the feature sparsity $S$ constant during training (blue solid line).
Thus, the core of DSFF mechanism is the learning of binary masks during the training of E2ENet. At initialization, each binary mask is initialized randomly, with the number of non-zero entries $\|\mathbf{M}^{j,i}\|_{0}$ equals $(1-S)\times C^{j,i}_{in}\times C^{j,i}_{out}$. Here, $S$ ($10<S<$ 1) is a hyperparameter called the feature sparsity level, which determines the percentage of feature map connections that are randomly inactivated. The activated connections can be updated throughout the course of training, while the feature sparsity level $S$ remains constant. This allows for both testing and training efficiency, as the exploitation of connections is maintained sparsely throughout training.
Every $\Delta T$ training epoch, the activated connections with lower importance are removed, while the same number of deactivated connections are randomly reactivated, as shown in Figure 4. The importance of activated connection is determined by the L1-norm of the corresponding kernel. That is, for $\mathbf{M}^{j,i}$, the importance score of connection between $c_{in}$-th input feature and $c_{out}$-th output is $\|\mathbf{\theta}^{j,i}_{c_{in},c_{out},:,:,:}\|_{1}$. Our intuition is simple: if the feature map connection is more important (larger effect on the output accuracy) than others, the L1-norm of the corresponding kernel should be larger. However, the importance score is suboptimal during training, as the kernels are gradually optimized. Randomly reactivating previous “obsolete” connections with re-initialization can avoid unrecoverable feature map abandonment and thoroughly explore the representative feature maps that contribute to the final performance.
The DSFF mechanism allows for the exploration and exploitation of multi-scaled sparse features through a novel sparse-to-sparse training method. Our method selects features at multiple scales within each layer, as opposed to relying solely on feature selection at the input layer (Sokar et al., 2022; Atashgahi et al., 2022). Additionally, it is a plastic approach to fuse features that can adapt to changing conditions during training, as opposed to static methods that rely on one-shot feature selection, as shown in the comparison of dynamic and static sparsity in Table 2 of (Mostafa & Wang, 2019). More details of the training processing are elaborated in Algorithm 1.
3.3 Restricted Depth-Shift in 3D Convolution
In 3D medical image segmentation, the 2D-based methods (such as the 2D nnUNet (Isensee et al., 2021)), which apply the 2D convolution to each slice of the 3D image, are computationally efficient but can not fully capture the relationships between slices. To overcome this limitation, we take inspiration from the temporal-shift (Lin et al., 2019) and 3D-shift (Fan et al., 2020) in efficient video action recognition, and the axial-shift (Lian et al., 2021) in efficient MLP architecture for vision. Our proposed E2ENet incorporates a depth-shift strategy in 3D convolution operations, which facilitates inter-slice information exchange and captures 3D spatial relationships while retaining the simplicity and computational efficiency of 2D convolution. Temporal-shift (Lin et al., 2019) requires selecting a Shift Proportion (the proportion of channels to conduct temporal shift). Axial-shift (Lian et al., 2021) and 3D-Shift optimize the learnable 3D spatiotemporal shift. We have made refinements to the channel shifting technique by shifting along the depth dimension and incorporating constraints on the shift size. This adaptation is thoughtfully designed to align with the distinctive needs of sparse models employed in medical image segmentation.
In our method, we employ a simple depth shift technique by shifting all channels while constraining the shift size to be either +1, 0 or -1, as shown in Figure 5. This choice is motivated using dynamic sparse feature fusion, where the feature maps contain sparse information. If only a portion of channels is selected for shifting or if the shift magnitude is too large, it can result in an insufficient representation of channels or an excessive representation of depth information, which can have a negative impact on the effectiveness of the shift operation (experimental results can be found in Table 5).
Figure 5: Illustration of restricted depth-shift in 3D Convolution of our E2ENet. The input features (left) are firstly split into 3 parts along the channel direction, and then shifted by $\{-1,0,1\}$ units along the depth dimension respectively (middle). After that, 3D CNNs with kernel size 1$\times$3$\times$3 is performed on the feature maps (middle) to generate the output features (right).
4 Experiments
In this section, we compare the performance of our E2ENet model to baseline methods on three datasets and report results in terms of both segmentation quality and efficiency. In addition, we will perform ablation studies to investigate the behavior of each component in the E2ENet model. To further analyze the performance of our model, we will present qualitative results by visualizing the predicted segmentations on sample images. We will also visualize the feature fusion ratios to gain insights into the importance of each feature in the segmentation process. Furthermore, we will discuss the relationship between organ volume and organ segment accuracy to understand the effect of organ size on the segmentation quality.
4.1 Datasets
@@ 4.2 Implementation Details
4.3 Evaluation Metrics
4.3.1 Mean Dice Similarity Coefficient
@@ 4.3.2 Number of Parameters
4.3.3 Float Point Operations
@@
4.3.4 Performance Trade-Off Score
@@
In most cases, we consider both segmentation accuracy and resource cost to be equally important, thus we set $\alpha_{1}=1$ and $\alpha_{2}=1/2$ in the following experiments. However, we also explore the impact of different choices of $\alpha_{1}$ and $\alpha_{2}$, as detailed in Section 4.6. The PT score serves as a valuable metric for evaluating the trade-offs between segmentation accuracy and efficiency.
Table 1: Quantitative comparisons (class-wise Dice (%) $\uparrow$, mDice(%$\uparrow$, Params(M)$\downarrow$, inference FLOPs(G)$\downarrow$, PT $\uparrow$ and mNSD(%$\uparrow$) of segmentation performance on the validation set of AMOS-CT dataset. Bold indicates the best and underline indicates the second best. Note: Spl: spleen, RKid: right kidney, LKid: left kidney, Gall: gallbladder, Eso: esophagus, Liv: liver, Sto: stomach, Aor: aorta IVC: inferior vena cava, Pan: pancreas, RAG: right adrenal gland, LAG: left adrenal gland, Duo: duodenum, Bla: bladder, Pro/Uth: prostate/uterus. The class-wise Dice, mDice and mNSD results of baselines, except for nnUNet, are collected from the (Ji et al., 2022). ${}^{{\dagger}}$ indicates the results without postprocessing that are collected from the AMOS website. ${}^{{\ddagger}}$ denotes the results with postprocessing that are reproduced by us. $*{}^{*}$ indicates the results with postprocessing.
@@
${}^{{\dagger}}$ 97.1 96.4 96.2 83.2 87.5 97.6 92.2 96.0 92.5 88.6 81.2 81.7 85.0 90.5 85.0 90.0 30.76 1067.89 1.30 82.1
${}^{{\ddagger}}$ 97.1 97.0 97.1 86.6 87.7 97.9 92.4 96.0 92.7 88.8 81.6 82.1 85.0 90.6 85.2 90.5 30.76 1067.89 1.31 83.0
$*{}^{*}$ (s=0.7) 97.1 96.9 97.1 86.0 87.6 97.9 92.3 95.7 92.3 89.0 81.5 82.4 84.9 90.3 83.8 90.3 11.23 969.32 1.54 82.7
$*{}^{*}$ (s=0.8) 97.1 96.9 97.0 85.2 87.5 97.9 92.3 95.7 92.3 89.0 81.3 82.1 84.6 90.1 84.8 90.3 9.44 778.74 1.65 82.5
$*{}^{*}$ (s=0.9) 96.7 96.9 97.0 84.2 87.0 97.7 92.2 95.6 92.0 88.6 81.0 81.8 84.0 89.9 83.8 89.9 7.64 492.29 1.89 81.8
@@
Table 2: Quantitative comparisons of segmentation performance on BTCV test set. Note: Spl: spleen, RKid: right kidney, LKid: left kidney, Gall: gallbladder, Eso: esophagus, Liv: liver, Sto: stomach, Aor: aorta IVC: inferior vena cava, Veins: portal and splenic veins, Pan: pancreas, AG: adrenal gland. The results (class-wise Dice and mDice) for these baselines are from (Hatamizadeh et al., 2022). ${}^{+}$ denotes that the training of ${}^{+}$ is without using any extra data outside the challenge. The results of ${}^{{\ddagger}}$, E2ENet and Hausdorff Distance (HD)$\downarrow$ of UNETR are from the standard leaderboard of BTCV challenge, while the results of nnUNet are from the free leaderboard.
@@
4.4 Performance Evaluation of E2ENet
In this section, we present the performance evaluation of our proposed E2ENet model in comparison to several state-of-the-art baseline models on three different challenges: AMOS-CT challenge, BTCV challenge, and BraTS challenge in MSD. Specifically, we compare the models in terms of Dice/mDice, Params, FLOPs, and PT score. The results demonstrate the effectiveness of E2ENet, which achieves comparable performance with fewer parameters and lower computational cost than many other models.
4.4.1 AMOS-CT Challenge
To fairly and comprehensively validate our method, we compare it to several state-of-the-art CNN-based models (e.g. nnUNet (Isensee et al., 2021), and Vnet (Milletari et al., 2016)) and transformer-based models (e.g. CoTr (Xie et al., 2021), nnFormer (Zhou et al., 2021), UNETR (Hatamizadeh et al., 2022), and Swin UNETR (Tang et al., 2022)). We recorded the class-wise Dice, mDice, Params, inference FLOPs, and PT score on the validation set [^4] The validation set includes 100 images and was the previous test set for the first stage of the challenge. In the training of E2ENet, we did not use the validation set and treated it as the test set. of the AMOS-CT challenge in Table 1. E2ENet with sparsity level $0.8$ achieves comparable performance with a mDice of $90.3\%$, while being four times smaller in size and requiring less computational cost in the testing phase, compared to the top-performing and lightweight model nnUNet. As the feature sparsity ($S$) of E2ENet increases, the number of model parameters and inference FLOPs can be decreased even further without significantly compromising segmentation performance. This indicates that there is potential to trade-off performance and efficiency by adjusting the sparsity of the model. We reported the results of normalized surface dice (NSD), which is the official segmentation metrics for AMOS challenge, to provide supplementary information on boundaries segmentation quliaty. We also presented the results of the test set in Table 8.
4.4.2 BTCV Challenge
4.4.3 BraTS Challenge in MSD
On the 5-fold validation of the training set, our E2ENet model demonstrates superior performance in terms of mDice compared to other state-of-the-art methods (UNETR (Hatamizadeh et al., 2022), DiNTS (He et al., 2021), nnUNet (Isensee et al., 2021) and UNet++ (Zhou et al., 2018)). Additionally, it is a computationally efficient network that is competitive among the baselines, as evidenced by its small model size and low inference FLOPs. Specifically, the E2ENet model with a feature sparsity level of $90\%$ has only $7.63$ M parameters, which is significantly smaller than other models yet still outperforms them in terms of mDice. Overall, the E2ENet model provides an optimal trade-off between performance and efficiency.
4.5 Ablation Studies
In this section, we investigate the impact of two factors on the performance of the E2ENet model: the DSFF mechanism and restricted depth-shift in 3D convolution. Specifically, we consider the following scenarios: (#1) w/ DSFF: DSFF mechanism is used to dynamically activate feature map, otherwise, all feature maps are activated during training; (#2) w/ shift: restricted depth/height/width shifting is applied on feature maps before the convolution operation, otherwise, the convolution is performed directly as a standard 3D convolution without shifting operation. We conduct ablation studies on the BraTS challenge and the AMOS-CT challenge. We opted not to include the BTCV challenge in our study due to its similarity to the AMOS-CT challenge, as both involve multi-organ segmentation in CT images. Furthermore, the AMOS-CT challenge is a larger and more comprehensive evaluation platform, allowing us to thoroughly assess the effectiveness of two factors. As for the BraTS challenge, it is specific to brain MRI image segmentation, which differs from the previous two challenges in focus and image modality.
Table 3: 5-fold cross-validation of segmentation performance on the training set of BraTS Challenge in the MSD. Note: ED, ET and NET denote edema, enhancing tumor and non-enhancing tumor, respectively.
Methods ED ET NET mDice Parmas FLOPs 1 PT score
DiNTS 80.2 61.1 77.6 73.0 / / /
UNet++ 80.5 62.5 79.2 74.1 58.38 3938.25 1.12
nnUNet 81.0 62.0 79.3 74.1 31.20 1076.62 1.35
E2ENet ($S=0.7$) 81.2 62.7 79.5 74.5 11.24 1067.06 1.57
E2ENet ($S=0.8$) 81.0 62.5 79.0 74.2 9.44 780.97 1.72
E2ENet ($S=0.9$) 80.9 62.5 79.4 74.3 7.63 494.52 2.00
  • 1
    The inference FLOPs are calculated based on patch sizes of 4 × 128 × 128 × 128. The number of parameters and inference FLOPs for DiNTS are not reported since calculating them can be time-consuming. This is due to the fact that the architecture for DiNTS is not readily available in this dataset and must be searched for using Neural Architecture Search (NAS).
Table 4: The ablation study of the effect of DSFF mechanism and restricted depth-shift in 3D convolution, evaluated through 5-fold cross-validation of segmentation performance on the training set of the BraTS Challenge in the MSD.
w/ DSFF shift kernel size ED ET NET mDice Params FLOPs
$(1\times 3\times 3)$ 80.4 62.4 79.0 73.9 23.89 3071.78
$(1\times 3\times 3)$ 81.0 62.3 79.0 74.1 23.89 3071.78
$(1\times 3\times 3)$ 80.3 62.5 79.0 73.9 11.24 1067.06
$(1\times 3\times 3)$ 81.2 62.7 79.5 74.5 11.24 1067.06
$(3\times 1\times 3)$ 80.3 61.7 78.7 73.6 11.24 1067.06
$(3\times 3\times 1)$ 80.5 61.9 78.7 73.7 11.24 1067.06
$(3\times 1\times 3)$ 81.1 62.4 78.9 74.1 11.24 1067.06
$(3\times 3\times 1)$ 81.0 62.3 79.4 74.2 11.24 1067.06
4.5.1 Effect of DSFF Mechanism
Table 4 shows that removing the DSFF mechanism (2nd row) caused the mDice to drop from $74.5\%$ to $74.1\%$ for the BraTS challenge, while also increasing the number of parameters and inference FLOPs by nearly 2-3 times. On the AMOS-CT challenge, Table 5 demonstrates that the E2ENet model with the DSFF mechanism achieved comparable performance with 3 times fewer parameters and inference FLOPs. These ablation study results highlight the effectiveness of the dynamic sparse feature fusion (DSFF) mechanism.
Moreover, as shown in Table 3, the E2ENet model with the DSFF mechanism outperforms other feature fusion architectures such as UNet++ and DiNTS in the BraTS challenge. Overall, the results of the BraTS and AMOS-CT challenges demonstrate that the DSFF mechanism is an effective and efficient feature fusion process with lower computational and memory overhead.
Table 5: The ablation study of the effect of DSFF mechanism and restricted depth-shift in 3D convolution on the validation set of AMOS-CT Challenge.
w/ DSFF $\Delta T$ (# iters) w/ shift shift size kernel size mDice Params FLOPs mNSD 1
/ $(-1,0,1)$ $(1\times 3\times 3)$ 90.2 23.90 3069.55 82.6
/ $(-1,0,1)$ $(1\times 3\times 3)$ 88.6 23.90 3069.55 78.6
1200 $(-1,0,1)$ $(1\times 3\times 3)$ 88.2 11.23 969.32 79.4
1200 $(-1,0,1)$ $(1\times 3\times 3)$ 90.1 11.23 969.32 82.3
1200 $(-2,0,2)$ $(1\times 3\times 3)$ 89.8 11.23 969.32 82.0
1200 $(-3,0,3)$ $(1\times 3\times 3)$ 89.7 11.23 969.32 81.6
1200 /// $(3\times 3\times 3)$ 90.1 52.54 4511.57 82.1
1200 /// $(3\times 3\times 3)$ 90.2 27.97 1778.55 82.5
600 $(-1,0,1)$ $(1\times 3\times 3)$ 90.0 11.23 969.32 82.2
1800 $(-1,0,1)$ $(1\times 3\times 3)$ 89.9 11.23 969.32 82.0
  • 1
    Inference FLOPs are calculated without including shift operations.
To investigate the impact of the DSFF module, we compared E2ENet with other multi-scale medical image segmentation methods on the AMOS-CT validation dataset. These methods include DeepLabv3 (Chen et al., 2017), CoTr (Xie et al., 2021), and MedFormer (Gao et al., 2022) (see Table 6). DeepLabv3 uses atrous convolution for capturing multi-scale context. CoTr integrates multi-scale features using attention, while MedFormer employs all-to-all attention for comprehensive multi-scale fusion, addressing both semantic and spatial aspects.
To further investigate the impact of the DSFF module, we compared E2ENet with other multi-scale medical image segmentation methods on the AMOS-CT validation dataset, including DeepLabv3 (Chen et al., 2017), CoTr(Xie et al., 2021), and MedFormer (Gao et al., 2022) (see Table 6). DeepLabv3 uses atrous convolution for multi-scale context capture. CoTr integrates multi-scale features using attention. MedFormer employs all-to-all attention for comprehensive multi-scale fusion, both semantically and spatially aware. We observed that, with a sparsity ratio of 0.9, E2ENet achieves performance comparable to DeepLabv3 while significantly reducing computational and memory costs, by more than 2x and 8x, respectively. E2ENet with a sparsity ratio of 0.8 matches MedFormer’s performance while significantly reducing computational and memory costs, by more than 2x and 4x, respectively. This demonstrates the benefits of multi-scale feature aggregation for medical image segmentation, with E2ENet’s DSFF module being much more efficient than other multi-scale methods. Finally, we also studied the impact of topology update frequency ($\Delta T$) in the DSFF module and observed that our algorithm is not sensitive to the hyperparameter $\Delta T.$
Table 6: The comparison with other multi-scale medical image segmentation methods on the AMOS-CT validation dataset. The mDice of MedFormer is collected from (Gao et al., 2022).
method mDice Params FLOPs
DeepLabv3 89.0 1546.25 74.68
CoTr 77.1 1510.53 41.87
MedFormer 90.1 2332.75 39.59
E2ENet (S=0.8) 90.0 1041.42 9.43
E2ENet (S=0.9) 89.0 680.25 7.63

### v2

@@ A.3 Algorithm
19:              $IS\leftarrow$ importance score ($L_{1}$ Norm of corresponding kernel) for activated each feature connection
@@
A.4 Datasets and Experiment Setup
@@ A.5 Implementation Details
A.6 The Architecture
The backbone generates a total of $L=6$ multi-scale feature levels, each with a specified number of channels: [c1, c2, c3, c4, c5, c6] = [48, 96, 192, 320, 320, 320]. At each level of feature generation, there are two convolution layers with a kernel size of (1, 3, 3), followed by instance normalization and the application of leaky ReLU activation. The down-sampling ratios for each level are as follows: ((1, 2, 2), (2, 2, 2), (2, 2, 2), (2, 2, 2), (2, 2, 2), (2, 2, 2), (2, 2, 2)).
A.7 Evaluation Metrics
A.7.1 Mean Dice Similarity Coefficient
@@ A.7.2 Number of Parameters
A.7.3 Float Point Operations
@@
A.7.4 Performance Trade-Off Score
@@
In most cases, we consider both segmentation accuracy and resource cost to be equally important, thus we set $\alpha_{1}=1$ and $\alpha_{2}=1/2$ in the following experiments. However, we also explore the impact of different choices of $\alpha_{1}$ and $\alpha_{2}$, as detailed in Section A.9. The PT score serves as a valuable metric for evaluating the trade-offs between segmentation accuracy and efficiency.
A.7.5 Discussion on Running Time
Given the relatively restricted support for sparse operations in current off-the-shelf commodity GPUs and TPUs without sparsity-aware accelerators, we did not attempt to achieve practical speedup during training. Instead, we chose to implement our models with binary masks in our work. The promising benefits of dynamic sparsity presented in this study have not yet translated into actual speedup. Accelerating training time will be a focus for our next work.
Although not the focus of our current work, it would be interesting for future work to examine the speedup results of sparse operation during training, using such specialized hardware accelerators. For example, at high unstructured sparsity levels, XNNPACK  [^7] https://github.com/google/XNNPACK. has already shown significant speedups over dense baselines on smartphone processors.
Table 9: Comparison of inference speeds between E2ENet and nnUNet.
Methods nnUNet E2ENet(s=0.8) E2ENet(s=0.9)
Latency(ms) 10.07 8.02 7.28
Throughput(items/sec) 99.19 124.60 137.13
speedup 1.0 $\times$ 1.26$\times$ 1.38$\times$
Although the support for unstructured sparsity on GPUs remains relatively limited, its practical relevance has been widely demonstrated on non-GPU hardware, such as CPUs and customized accelerators. For example, FPGA accelerators have demonstrated significant acceleration and energy efficiency for unstructured sparse RNNs, outperforming commercial CPUs and GPUs by effectively leveraging the FPGA’s embedded resources. Additionally, DeepSparse  [^8] https://github.com/neuralmagic/deepsparse. has shown impressive results in deploying large-scale, BERT-level sparse models on modern Intel CPUs. This approach achieved a 10× compression in model size with under a 1% accuracy loss, a 10× increase in CPU inference speed with less than a 2% accuracy reduction, and an impressive 29× CPU inference speedup with a maximum accuracy drop of 7.5% (Liu and Wang, 2023).
Inspired by these advancements, we adopted an approach based on DeepSparse. As shown in Table 9, we conducted experiments with patches of images sized 32×32×32 as input, comparing the CPU wall-clock timings for online inference between our proposed E2ENet and nnUNet on an Intel Xeon Platinum 8360Y CPU with 18 cores. We acknowledge that, while our proposed models with sparsity do achieve speedups in practical inference, they are not as pronounced as those observed with BERT-level sparse models. This is primarily due to the nature of segmentation and 3D convolution operations. However, this presents a promising avenue for our future work.
A.8 More Experimental Results
A.8.1 Class-wise Dice of AMOS-CT
Table 10: Quantitative comparisons (class-wise Dice (%) $\uparrow$, mDice(%$\uparrow$, Params(M)$\downarrow$, inference FLOPs(G)$\downarrow$, PT $\uparrow$ and mNSD(%$\uparrow$) of segmentation performance on the validation set of AMOS-CT dataset. Bold indicates the best and underline indicates the second best. Note: Spl: spleen, RKid: right kidney, LKid: left kidney, Gall: gallbladder, Eso: esophagus, Liv: liver, Sto: stomach, Aor: aorta IVC: inferior vena cava, Pan: pancreas, RAG: right adrenal gland, LAG: left adrenal gland, Duo: duodenum, Bla: bladder, Pro/Uth: prostate/uterus. The class-wise Dice, mDice and mNSD results of baselines, except for nnUNet, are collected from the (Ji et al., 2022). indicates the results without postprocessing that are collected from the AMOS website. denotes the results with postprocessing that are reproduced by us. indicates the results with postprocessing.
@@ A.8.1 Class-wise Dice of AMOS-CT
nnUNet 97.1 96.4 96.2 83.2 87.5 97.6 92.2 96.0 92.5 88.6 81.2 81.7 85.0 90.5 85.0 90.0 30.76 1067.89 1.30 82.1
nnUNet 97.1 97.0 97.1 86.6 87.7 97.9 92.4 96.0 92.7 88.8 81.6 82.1 85.0 90.6 85.2 90.5 30.76 1067.89 1.31 83.0
E2ENet (s=0.7) 97.1 96.9 97.1 86.0 87.6 97.9 92.3 95.7 92.3 89.0 81.5 82.4 84.9 90.3 83.8 90.3 11.23 969.32 1.54 82.7
E2ENet (s=0.8) 97.1 96.9 97.0 85.2 87.5 97.9 92.3 95.7 92.3 89.0 81.3 82.1 84.6 90.1 84.8 90.3 9.44 778.74 1.65 82.5
E2ENet (s=0.9) 96.7 96.9 97.0 84.2 87.0 97.7 92.2 95.6 92.0 88.6 81.0 81.8 84.0 89.9 83.8 89.9 7.64 492.29 1.89 81.8
@@
A.8.2 BTCV Challenge
Table 11: Quantitative comparisons of segmentation performance on BTCV test set. Note: Spl: spleen, RKid: right kidney, LKid: left kidney, Gall: gallbladder, Eso: esophagus, Liv: liver, Sto: stomach, Aor: aorta IVC: inferior vena cava, Veins: portal and splenic veins, Pan: pancreas, AG: adrenal gland. The results (class-wise Dice and mDice) for these baselines are from (Hatamizadeh et al., 2022). + denotes that the training of UNETR+ is without using any extra data outside the challenge. The results of nnUNet, E2ENet and Hausdorff Distance (HD)$\downarrow$ of UNETR are from the standard leaderboard of BTCV challenge, while the results of nnUNet are from the free leaderboard.
@@ A.8.2 BTCV Challenge
A.8.3 Statistical Significance of Designed Modules